# Induced Berry Connection and Photonic Spin Hall Effect in Optical Dirac Theory

### v1

@@ III EQUATION OF MOTION
$\hat{z}_{i}=(\hat{r}_{i}+i\hat{p}_{i})/\sqrt{2},\quad\hat{z}^{*}_{i}=(\hat{r}_{i}-i\hat{p}_{i})/\sqrt{2}$ (29)
@@
$\dot{r}_{i}=\frac{\partial\mathcal{H}_{H}}{\partial p_{i}}-\frac{\partial\mathcal{H}_{A}}{\partial r_{i}}\quad\dot{p}_{i}=-\frac{\partial\mathcal{H}_{H}}{\partial r_{i}}-\frac{\partial\mathcal{H}_{A}}{\partial p_{i}}$ (35)
In the Hermitian part of the Hamiltonian (6), the TAM-orbit coupling term $\mathbf{\Omega}\cdot\mathbf{J}$ consists of two distinct contributions, one arising from the photon’s spin $\sigma$ and another from its topological charge $l$, both of which will lead to a unified gauge structure in the momentum space. In the form of path integrals, they play a critical role to parallel transport state vectors $|\mathbf{e}_{\mathbf{k}}\rangle$ to adjacent points $|\mathbf{e}_{\mathbf{k+\vartriangle k}}\rangle$ as [30]:
$\langle\mathbf{e}_{\mathbf{k+\vartriangle k}}|\mathbf{e}_{\mathbf{k}}\rangle=e^{-i\mathbf{A}_{k}(\mathbf{k})\cdot\dot{\mathbf{k}}}$ (36)
where $\mathbf{A}_{k}$ represents the Berry connection in momentum space. It is related to torsion $\tau=-\mathbf{A}\dot{\mathbf{k}}$ in momentum space representation [31]. The corresponding Berry curvature is $\mathbf{F}=\partial_{\mathbf{k}}\times\mathbf{A}_{k}$, exhibiting the structure of a monopole [25]:
$\mathbf{F}(k)=(l+\sigma)\frac{\mathbf{k}}{k^{3}}$ (37)
which will give rise to a non-integral topological phase - the Berry phase - as given in the equation (17).
An additional contribution comes from the OAM of the wave packet, which arises from the spatial arrangement of photons relative to their center of mass. This OAM changes along with the helical optical path that gives rise to the rotation of the reference frame, a phenomenon analogous to a "Coriolis force".
@@
The first term in the right-hand side of Eqn.(38) describes the Lamor precession of a photon characterized by the angular velocity $-\mathbf{\Omega}$ along the helical optical path. The second term introduces a "Lorentz force" in momentum space, giving rise to the spin (or orbital) Hall effect of light [25, 31]. The first term in the right-hand side of the motion equation Eqn.(39) features a "Coriolis force" term, $\mathbf{\Omega}\times\mathbf{k}$ [32]. The directional derivative of $\Gamma_{h}$ and $\Gamma_{a}$, either in real space $\mathbf{r}$ or in momentum space $\mathbf{k}$, are analyzed by decomposing the first order of $1/k_{z}$ in Hamiltonians $P_{0}\gamma_{0}+P_{\perp}\cdot\gamma_{\perp}$ into Hermitian and non-Hermitian components $P_{i}=P^{h}_{i}-iP^{a}_{i}$. Accordingly, $\Gamma_{h}=\langle\mathbf{\Psi}|P^{h}_{\perp}\cdot\gamma_{\perp}+P^{h}_{0}\gamma_{0}|\mathbf{\Psi}\rangle$ represents the Hermitian part, while $\Gamma_{a}=\langle\mathbf{\Psi}|P^{a}_{\perp}\cdot\gamma_{\perp}+P^{a}_{0}\gamma_{0}|\mathbf{\Psi}\rangle$ corresponds to the anti-Hermitian part. These components account for non-Abelian contributions, as discussed in [33]. At last, the terms $\mathbf{v}_{s}$ and $\mathbf{f}_{s}$ result from the contribution of the paraxial field, explicitly, given by
@@
The expectation value of physical observables is evaluated over quantum states characterized by the wave function, $\mathbf{\Psi}$, which encompasses contributions from both positive and negative photon energy states. Consequently, classical physical quantities derived from the wave function must also account for the contributions of antiphotons. For example, an antiphoton can be understood as a photon traveling in the opposite direction. For instance, the mean value of the momentum operator in a given state $\mathbf{\Psi}$ can be written as:
@@
where $\mathbf{F}_{\mathbf{k}}$ and $\mathbf{F}_{\mathbf{r}}$ represent field-strength tensors associated with the gauge fields. It is evident that the Lorentz force described in equation 38 results from this non-commutative relation, where $\mathbf{F}_{\mathbf{k}}$ is defined in terms of equation (37). Thus, the Hall effect terms can also be interpreted within the framework of non-commutative geometry, as discussed in [35, 36, 37].
IV EVOLUTION OF SPIN AND ORBIT ANGULAR MOMENTUM
@@
where $L^{ext}_{z}$ includes contributions from the transverse shift, linked to both the SHE[40, 41, 42] and the OHE [43, 44, 45, 46].
@@
V INTRINSIC SPIN-ORBIT INTERACTIONS
@@
$\mathbf{E}=A_{0}\left(\mathbf{e}_{x}-i\frac{\kappa}{k_{z}}\mathbf{e}_{z}\right)\exp\left(ik_{z}z-\kappa x\right)$ (66)
@@
$\mathbf{k}=k\mathbf{e}_{z}+i\mathbf{\kappa}\mathbf{e}_{x}$ (67)
Here, $\kappa$ represents the decay constant in the transverse x-direction, and $k_{z}$ is the longitudinal component along the z-direction. Due to the transverse condition $\mathbf{\nabla}\cdot\mathbf{E}=\mathbf{k}\cdot\mathbf{E}=0$, we obtain the imaginary component $E_{z}$ as:
$E_{z}=-i\frac{\kappa}{k_{z}}E_{x}$ (68)
@@
$\mathbf{S}\varpropto Im(\mathbf{E}^{*}\times\mathbf{E})$ (69)
Accordingly, for the electric field given by Eq.(66), the transverse spin density can be readily determined as, $s_{y}\propto\kappa/k_{z}$ [47]. Equivalently, the spin density $s_{y}$ can also be expressed as the expectation value of the spin operator in the wave function representation [59, 60, 61].
@@
Again, we have $s_{y}\propto\kappa/k_{z}$ using Eq.(68), .
@@
$\mathbf{S}_{\perp}=\frac{Re\mathbf{k}\times Im\mathbf{k}}{(Re\mathbf{k})^{2}}$ (74)
@@
VI SUMMARY and CONCLUDING REMARKS
@@ Appendix A OPTICAL DIRAC EQUATION IN HELICAL COORDINATES
$\begin{array}[]{ll}i\displaystyle{\frac{\partial\mathcal{D_{\perp}}}{\partial t}}=(H_{0}^{B}-{\bf H}^{B}\cdot{\bm{\sigma}}_{\perp}+H_{3}^{B}\sigma_{3})i{\sigma}_{3}{\mathcal{B}}_{\perp}\\ \\ i\displaystyle{\frac{\partial i\sigma_{3}\mathcal{B}_{\perp}}{\partial t}}=(H_{0}^{D}+{\bf H}^{D}\cdot{\bm{\sigma}}_{\perp}+H_{3}^{D}\sigma_{3}){\mathcal{D}}_{\perp}\end{array}$ (106)
where the superscript $D$ and $B$ denote the corresponding scalar quantities and matrices related to the operators ×$\nabla\times\epsilon^{-1}$ and ×$\nabla\times\mu^{-1}$. Introducing the two-vector wavefunctions of photons in the transverse helicity space,
@@
we further define a four-vector ${\bf\Psi}=({\bf\Psi}_{+},{\bf\Psi}_{-})^{T}$. Based on these definitions, the Maxwell equation Eqn.(106) can be converted into a compact form of the Dirac equation with the chiral extension,
@@
in which the effective masses and momentums are
@@
${\bf H}_{\perp}^{\pm}$ $=$ $\displaystyle{\frac{1}{2}}({\bf H}^{D}\pm{\bf H^{B}})$ (111)
@@
Now, we apply the above procedure to obtain the optical Dirac equation in the helical coordinate frame specified by the spatial metric $\Gamma=\{\gamma_{\alpha\beta}\}$ as given in Eq.(2). Under the unitary transformation $\hat{U}$, we have
@@
Using the transverse components of $\mathbf{D}$, we obtain the transverse components under the Hamiltonian $\hat{H}$
@@
$\hat{H}_{\perp}=\sigma_{3}\Bigl{[}\hat{k}_{z}-\tau(\hat{L}_{z}+\sigma_{3})+\frac{1}{\hat{k}_{z}}(P_{I}+P_{\perp}\cdot\sigma_{\perp}+P_{3}\cdot\sigma_{3})\Bigr{]}$ (117)
@@
$P_{\pm}$ $=g_{s}k^{2}_{\pm}+2i(\tau k_{z}r_{\pm}+g_{\pm})k_{\pm}$ (119)
$P_{3}$ $=({\bf g}_{\perp}\times{\bf k}_{\perp})_{z}$ (120)
where $g_{s}=\gamma_{00}$, $\mathbf{g}_{\perp}=\frac{1}{2}\nabla_{\perp}g_{s}$. The first term in $P_{0}$ is the conventional transverse-momentum correction in the paraxial approximation, the second gives the dipole interaction. Ladder operators $\sigma_{pm}$ serve to flip spin by two angular momentum units, facilitating the first term in $P_{\pm}$ of Eq.(119), which represents the SAM-IOAM conversion of photons. The second term explains how SAM and IOAM of photons interact with the spin-1 dipole mode of the optical medium. Each term ensures the conservation of angular momentum individually. The $P_{3}$ term is formally a spin-orbit coupling ${\bf L}_{g}\cdot{\bm{\sigma}}$ with ${\bf L}_{g}={\bf g}_{\perp}\times{\bf k}_{\perp}$. It is noted that the $1/k_{z}$ terms are from the paraxial approximation and are smaller than the zeroth leading term by a factor of $1/wk_{z}$ ($w$ is the waist size of paraxial optical fields).
While considering a gyrotropic medium, there exist antisymmetric components in the dielectric tensor, such as:
$\gamma_{12}=-\gamma_{21}=-i\Lambda$ (121)
which, according to Eq.(102), leads to an additional rotation $\Lambda k_{z}$. Thus, the results regarding the helical optical path in this study can be extended to gyrotropic media simply by the replacement of $\Omega\rightarrow\Omega+\Lambda k_{z}$.

### v2

@@ IV EQUATION OF MOTION
$\hat{z}_{i}=(\hat{r}_{i}+i\hat{k}_{i})/\sqrt{2},\quad\hat{z}^{*}_{i}=(\hat{r}_{i}-i\hat{k}_{i})/\sqrt{2}$ (30)
@@
$\dot{r}_{i}=\frac{\partial\mathcal{H}_{H}}{\partial k_{i}}-\frac{\partial\mathcal{H}_{A}}{\partial r_{i}}\quad\dot{k_{i}}=-\frac{\partial\mathcal{H}_{H}}{\partial r_{i}}-\frac{\partial\mathcal{H}_{A}}{\partial k_{i}}$ (36)
In the Hermitian part of the Hamiltonian (6), the TAM-orbit coupling term $\mathbf{\Omega}\cdot\mathbf{J}$ consists of two distinct contributions, one arising from the photon’s spin $\sigma$ and another from its topological charge $l$, both of which will lead to a unified gauge structure in the momentum space.
Next, we proceed to derive the equations of motion while incorporating the effects of geometric phases. To account for these effects, we adopt the interaction representation, where both quantum state vectors and operators exhibit explicit time evolution. In this framework, the evolution of the state vector is governed by the Berry connection. The system’s Hamiltonian is conventionally partitioned as
$H=H_{0}+H_{I}$ (37)
where $H_{0}$ governs the operator dynamics while $H_{I}$ dictates the state vector evolution. The development of the state vector $|\psi(s)\rangle$ along the trajectory of light is governed by the path-ordered evolution operator:
$|\psi(s)\rangle=U_{I}(s,s_{0})|\psi(s_{0})\rangle$ (38)
which satisfies the fundamental dynamical equation:
$i\frac{d}{ds}U_{I}(s,s_{0})=H_{I}(s)U_{I}(s,s_{0})$ (39)
The formal solution admits a path-ordered exponential representation:
[…]
where $\mathcal{P}$ enforces path-ordering along the optical trajectory. In the path integral representation, the eigenvalues of the evolution operator correspond to the sum of the geometric phase and the dynamical phase [36]. Within the interaction picture, the operator $U_{I}$ governs the geometric contribution to state evolution. For infinitesimal momentum variations $\mathbf{k}\rightarrow\mathbf{k}+\Delta\mathbf{k}$, the overlap between adjacent eigenstates $|\psi_{k}\rangle$ in momentum-space satisfies
$\langle\psi_{\mathbf{k}+\Delta\mathbf{k}}|\psi_{\mathbf{k}}\rangle=1-i\langle\psi_{\mathbf{k}}|i\Delta\mathbf{k}\cdot\partial_{\mathbf{k}}|\psi_{\mathbf{k}}\rangle\approx e^{-i\mathbf{A}_{k}(\mathbf{k})\cdot\Delta\mathbf{k}}$ (41)
where $\mathbf{A}_{k}=\langle\psi_{\mathbf{k}}|i\partial_{\mathbf{k}}|\psi_{\mathbf{k}}\rangle$ denotes the Berry connection, governing eigenstate parallel transport in parameter space. By employing a path integral formulation, the contribution of (41) to the geometric phase can be expressed as[36]:
[…]
where $\Gamma$ represents a closed trajectory in momentum space. Consequently, the term $A_{\mathbf{k}}\cdot\dot{\mathbf{k}}$ naturally emerges as the interaction Hamiltonian $H_{I}$, which governs the evolution of the state vector in momentum space. In our system, this geometric evolution is characterized by the Berry curvature tensor[29]:
$\mathbf{F}(k)=\partial_{\mathbf{k}}\times\mathbf{A}_{k}=(l+\sigma)\frac{\mathbf{k}}{k^{3}}$ (43)
which exhibits a characteristic monopole structure in momentum space and give rise to a non-integral topological phase - the Berry phase - as given in the equation (16).
@@
The second term in the right-hand side of Eqn.(44) describes the Lamor precession of a photon characterized by the angular velocity $-\mathbf{\Omega}$ along the helical optical path. The third term introduces a "Lorentz force" in momentum space, giving rise to the spin (or orbital) Hall effect of light [8, 29, 11]. The first term in the right-hand side of the motion equation Eqn.(45) features a "Coriolis force" term, $\mathbf{\Omega}\times\mathbf{k}$. This additional contribution comes from the OAM of the wave packet, which arises from the spatial arrangement of photons relative to their center of mass. The directional derivative of $\Gamma_{h}$ and $\Gamma_{a}$, either in real space $\mathbf{r}$ or in momentum space $\mathbf{k}$, are analyzed by decomposing the first order of $1/k_{z}$ in Hamiltonians $P_{0}\gamma_{0}+P_{\perp}\cdot\gamma_{\perp}$ into Hermitian and non-Hermitian components $P_{i}=P^{h}_{i}-iP^{a}_{i}$. Accordingly, $\Gamma_{h}=\langle\mathbf{\Psi}|P^{h}_{\perp}\cdot\gamma_{\perp}+P^{h}_{0}\gamma_{0}|\mathbf{\Psi}\rangle$ represents the Hermitian part, while $\Gamma_{a}=\langle\mathbf{\Psi}|P^{a}_{\perp}\cdot\gamma_{\perp}+P^{a}_{0}\gamma_{0}|\mathbf{\Psi}\rangle$ corresponds to the anti-Hermitian part. These components account for non-Abelian contributions, as discussed in [37]. At last, the terms $\mathbf{v}_{s}$ and $\mathbf{f}_{s}$ result from the contribution of the paraxial field, explicitly, given by
@@
The expectation value of physical observables is evaluated over quantum states characterized by the wave function, $\mathbf{\Psi}$, which encompasses contributions from both positive and negative photon energy states. Consequently, classical physical quantities derived from the wave function must also account for the contributions of antiphotons. For instance, the mean value of the momentum operator in a given state $\mathbf{\Psi}$ can be written as:
@@
where $\mathbf{F}_{\mathbf{k}}$ and $\mathbf{F}_{\mathbf{r}}$ represent field-strength tensors associated with the gauge fields. It is evident that the Lorentz force described in Eq.(44) results from this non-commutative relation, where $\mathbf{F}_{\mathbf{k}}$ is defined in terms of Eq.(43). Thus, the Hall effect terms can also be interpreted within the framework of non-commutative geometry, as discussed in [39, 40, 41].
V EVOLUTION OF SPIN AND ORBIT ANGULAR MOMENTUM
@@
where $L^{ext}_{z}$ includes contributions from the transverse shift, linked to both the SHE and the OHE [8, 44, 45, 12, 46, 47, 10].
@@
VI INTRINSIC SPIN-ORBIT INTERACTIONS
@@
$\mathbf{E}=A_{0}\left(\mathbf{e}_{x}-i\frac{\kappa_{x}}{k_{z}}\mathbf{e}_{z}\right)\exp\left(ik_{z}z-\kappa_{x}x\right)$ (72)
@@
$\mathbf{k}=k\mathbf{e}_{z}+i\kappa_{x}\mathbf{e}_{x}$ (73)
Here, $\kappa_{x}$ represents the decay constant in the transverse x-direction, and $k_{z}$ is the longitudinal component along the z-direction. Due to the transverse condition $\mathbf{\nabla}\cdot\mathbf{E}=\mathbf{k}\cdot\mathbf{E}=0$, we obtain the imaginary component $E_{z}$ as:
$E_{z}=-i\frac{\kappa_{x}}{k_{z}}E_{x}$ (74)
@@
$\mathbf{S}\varpropto{\text{Im}}(\mathbf{E}^{*}\times\mathbf{E})$ (75)
Accordingly, for the electric field given by Eq.(72), the transverse spin density can be readily determined as, $s_{y}\propto\kappa_{redx}/k_{z}$ [7]. Equivalently, the spin density $s_{y}$ can also be expressed as the expectation value of the spin operator in the wave function representation [59, 60, 61].
@@
Again, we have $s_{y}\propto\kappa_{x}/k_{z}$ using Eq.(74), .
@@
$\mathbf{S}_{\perp}=\frac{\text{Re}\mathbf{k}\times\text{Im}\mathbf{k}}{(\text{Re}\mathbf{k})^{2}}$ (80)
@@
VII SUMMARY and CONCLUDING REMARKS
@@
It is necessary to emphasize here that, to demonstrate SOI of light, we examine the case of light propagating through a uniform helical optical fiber. While this represents a limiting case, the underlying conclusions can be generalized to more complex configurations for two primary reasons - Firstly, SOI and the spin-Hall effect of light are typically treated within the framework of semi-classical theory, where the notion of a classical trajectory is utilized to describe subtle beam shifts. Basically, a ray system associated with electromagnetic waves is determined by a Hamiltonian structure induced from the dispersion relation. In our approach, this ray system is reformulated in terms of a local curvilinear coordinate system adapted to the light’s propagation path. Secondly, the eikonal approximation (i.e., geometric optics approximation) underpins most semi-classical treatments. On the other hand, the Frenet–Serret coordinate system provides a general mathematical framework for describing spatial curves in three dimensions. Under the eikonal approximation, it is reasonable to approximate the light’s trajectory locally as a piecewise-smooth curve with constant curvature and torsion, which is actually equivalent to the adiabatic transportation of a quantum system.
@@ Appendix A OPTICAL DIRAC EQUATION IN HELICAL COORDINATES
$\begin{array}[]{ll}i\displaystyle{\frac{\partial\mathcal{D_{\perp}}}{\partial t}}=(H_{0}^{B}-{\bf H}^{B}_{\perp}\cdot{\bm{\sigma}}_{\perp}+H_{3}^{B}\sigma_{3})i{\sigma}_{3}{\mathcal{B}}_{\perp}\\ \\ i\displaystyle{\frac{\partial i\sigma_{3}\mathcal{B}_{\perp}}{\partial t}}=(H_{0}^{D}+{\bf H}^{D}_{\perp}\cdot{\bm{\sigma}}_{\perp}+H_{3}^{D}\sigma_{3}){\mathcal{D}}_{\perp}\end{array}$ (112)
where $\mathcal{D}_{\perp}$ and $\mathcal{B}_{\perp}$ are the transverse components of electromagnetic vectors ${\mathbf{D}}$ and ${\mathbf{B}}$ in the helicity space, the superscript $D$ and $B$ denote the corresponding scalar quantities and matrices related to the operators ×$\nabla\times\epsilon^{-1}$ and ×$\nabla\times\mu^{-1}$, respectively. Introducing the two-vector wavefunctions of photons in the transverse helicity space,
@@
we further combine them to form a four-vector ${\bf\Psi}{{}_{\perp}}=({\bf\Psi}_{+},{\bf\Psi}_{-})^{T}$. Based on these definitions, the Maxwell equation Eqn.(112) can be converted into a compact form of the Dirac equation with the chiral extension,
@@
in which $\gamma$ are the Dirac representation of gamma matrices, the effective masses and momenta are
@@
${\bf H}_{\perp}^{\pm}$ $=$ $\displaystyle{\frac{1}{2}}({\bf H}^{D}_{\perp}\pm{\bf H}^{B}_{\perp})$ (117)
@@
Now, we apply the above procedure to obtain the optical Dirac equation in the helical coordinate frame specified by the spatial metric $\Gamma=\{\widetilde{\gamma}_{\alpha\beta}\}$ as given in Eq.(2). Under the unitary transformation $\hat{U}$, we have
@@
Acting on the transverse components of $\mathbf{D}$, we obtain the transverse components under the Hamiltonian $\hat{H}$
@@
$\hat{H}_{\perp}=\sigma_{3}\Bigl{[}\hat{k}_{z}-\tau(\hat{L}_{z}+\sigma_{3})+\frac{1}{\hat{k}_{z}}(P_{0}+{\mathbf{P}}_{\perp}\cdot{\bm{\sigma}}_{\perp})\Bigr{]}$ (123)
@@
$P_{0}=P_{I}+P_{3}\sigma_{3}$ (124)
${\mathbf{P}}_{\perp}\cdot{\bm{\sigma}}_{\perp}=P_{+}\sigma_{+}+P_{-}\sigma_{-}$
with
@@
$P_{\pm}=g_{s}k^{2}_{\pm}+2i(\tau k_{z}r_{\pm}+g_{\pm})k_{\pm}$
$P_{3}=({\bf g}_{\perp}\times{\bf k}_{\perp})_{z}$
where $g_{s}=\widetilde{\gamma}_{00}$, $\mathbf{g}_{\perp}=\frac{1}{2}\nabla_{\perp}g_{s}$. The first term in $P_{0}$ is the conventional transverse-momentum correction in the paraxial approximation, the second gives the dipole interaction. Ladder operators $\sigma_{\pm}$ serve to flip spin by two angular momentum units, facilitating the first term in $P_{\pm}$ of Eq.(125), which represents the SAM-IOAM conversion of photons. The second term explains how SAM and IOAM of photons interact with the spin-1 dipole mode of the optical medium. Each term ensures the conservation of angular momentum individually. The $P_{3}$ term is formally a spin-orbit coupling ${\bf L}_{g}\cdot{\bm{\sigma}}$ with ${\bf L}_{g}={\bf g}_{\perp}\times{\bf k}_{\perp}$. It is noted that the $1/k_{z}$ terms are from the paraxial approximation and are smaller than the zeroth leading term by a factor of $1/wk_{z}$ ($w$ is the waist size of paraxial optical fields).
While considering a gyrotropic medium, there exist antisymmetric components in the dielectric tensor, such as $\widetilde{\gamma}_{12}=-\widetilde{\gamma}_{21}=-i\Lambda$, which, according to Eq.(108), leads to an additional rotation $\Lambda k_{z}$. Thus, the results regarding the helical optical path in this study can be extended to gyrotropic media simply by the replacement of $\Omega\rightarrow\Omega+\Lambda k_{z}$.